# Mitigating distribution shift in machine learning-augmented hybrid simulation

### v1

@@ 3.2 Algorithm in the general form
0:  $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$, resolved model, penalty strength $\lambda$.
1:  Learn a parameterized model which encode the structure of training data, i.e. $F_{\eta}(\mathbf{u})\geq 0,F_{\eta}(\mathbf{u}_{k})=0.$
@@
6:     Calculate the state variable after one-step iteration, i.e. $\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+\Delta t\mathcal{L}(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k},t_{k})$.
7:     Form the loss $l(\theta)=\mathbb{E}\left[\left\|\mathbf{y}_{k}-\widehat{\mathbf{y}}_{k}\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{k})\right)^{T}\mathcal{L}(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k}),t_{k})\right)^{2}\right]$.
@@
Comparing to literature of distributionally robust optimization (DRO) [40] and numerical analysis of dynamical systems [48, 47], our work takes advantages from both. First, the tangent-space regularized estimator belongs to the family of DRO in the sense that it minimizes the data-driven module of the unresolved part over some perturbations of the data distribution. However, unlike the original DRO which considers all possible perturbations of the data distribution in its neighborhood under some metric, e.g. Wasserstein metric  [3], our regularization focus on those perturbations caused by the resolved components of the dynamics. These facilitate us to take advantage of the partial knowledge of the dynamics. As a result, the regularizer we proposed is more problem-specific Moreover, from the perspective of DRO, our method takes into consideration the behavior of estimator under certain perturbation normal to the data manifold, which is much more tractable than guarding against arbitrary perturbations in distribution space. In Section 4, we will implement other benchmarks with general regularizations and compare with our algorithm to illustrate the benefit of combining resolved model information into regularization.
Furthermore, we can understand our method through the lens of the stability in numerical analysis of differential equations. The classical result that the least squares estimator converges to the ground truth in the limit of large datasets can be understood as a consistency statement. Here, we show that this is insufficient, and in order to ensure convergence of the simulated trajectories one needs to also promote stability, and our proposed regularizer serves precisely this role. This regularization approach should be contrasted with recent works [54, 7, 25, 16] on learning dynamical models from data which build stability by specifying the model architecture.
@@ 3.3 Interpretation and analysis of the algorithm
Let us now quantify in the linear case, the gains of our algorithm over the ordinary least squares estimator. We first calculates the exact formula for two estimators. Due to the low rank structure, OLS algorithm allows infinitely many solutions. Throughout this section, we make the common choice of the minimum 2-norm solution.
@@
Consider the linear dynamics Eq. 12, the training data of the variable $\mathbf{u}$ is arranged into a data matrix $\mathbf{U}\in\mathbb{R}^{m\times N}$. Moreover, we assume the column space of $\mathbf{U}$ is contained in a subspace $V\subset\mathbb{R}^{m}$ with associated projection operator $P_{V}$. Then, the OLS estimator and tangent-space regularized estimator for the unresolved component $Y=C^{*}\mathbf{U}+\epsilon$ have the following form:
$\widehat{C}_{OLS}=$ $\ C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger},$ (28)
@@
$\widehat{C}_{TR}=$ $\ (\mathbf{I}+\lambda B^{T}P_{V^{\perp}}B)^{-1}(C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger}-\lambda B^{T}P_{V^{\perp}}AP_{V}).$
@@
Specifically, in the noiseless scenarios, $\widehat{C}_{OLS}$ recovers $C^{*}$ on the subspace $V$ but vanishes on its orthogonal complement. In order to interpret our estimator, let us consider a simpler case with $B=\mathbf{I},A=\mathbf{0}$. This corresponds to the dynamics
$\partial_{t}\mathbf{u}=C^{*}\mathbf{u},$ (29)
and the regularized estimator is given by
@@
$\widehat{C}_{TR}=(\mathbf{I}+\lambda P_{V^{\perp}})^{-1}(C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger})=(\mathbf{I}+\lambda P_{V^{\perp}})^{-1}\widehat{C}_{OLS}.$ (30)
@@
Therefore, our estimator performs a weighted least squares regression which penalizes the direction perpendicular to the data subspace $V$, as controlled by $\lambda$. As the error along this normal direction is the main cause of the distribution shift, it is sensible to reduce it more than the error along the tangent direction, which cause deviations inside the manifold. This shows that the tangent-space regularized estimator does in fact mitigate distribution shift.
@@
$l_{TR}(\widehat{C})=\mathbb{E}\left(\left\|(C^{*}-\widehat{C})\mathbf{u}\right\|_{2}^{2}+\lambda\left\|P_{V^{\perp}}(A+B\widehat{C})\mathbf{u}\right\|_{2}^{2}\right)$ (31)
@@
$Q(r,T)=\frac{e^{rT}-1}{r},\quad r,T>0.$ (32)
@@
Under the same setting as Proposition 3.1, assume both the OLS estimator and TR estimator $\widehat{C}_{OLS},\widehat{C}_{TR}$ are optimized to have error bounded by $\delta$ so that $l_{OLS}(\widehat{C}_{OLS}),l_{TR}(\widehat{C}_{TR})<\delta$ in Eq. 15 and Eq. 31. Suppose the true trajectory is simulated from an initial condition $\mathbf{u}(0)\in V$ which follows the same distribution of the training data. Then, the errors of OLS and our algorithm are bounded respectively by
@@
$\mathbb{E}\left\|\widehat{\mathbf{u}}_{TR}(T)-\mathbf{u}(T)\right\|^{2}\leq$ $\delta Q(\left\|A+B\widehat{C}_{TR}P_{V}\right\|^{2}+3,T)$ (33)
×$\times\left(\left\|B\right\|^{2}+\frac{Q(\left\|P_{V^{\perp}}(A+B\widehat{C}_{TR})\right\|^{2}+2,T)}{\lambda}\right),$
@@
$\mathbb{E}\left\|\widehat{\mathbf{u}}_{OLS}(T)-\mathbf{u}(T)\right\|^{2}$ $\leq\delta Q(\left\|A+B\widehat{C}_{OLS}P_{V}\right\|^{2}+\left\|B\widehat{C}_{OLS}\right\|^{2}+3,T)\left\|B\right\|^{2}.$
The proof idea is illustrated in deriving Eq. 19 and we leave step-by-step derivations in the supplementary materials Section 3. To understand this result, let us focus on the exponential part of two error bounds as this will dominate the error when the simulation time $t$ is large. The exponent of our algorithm is bounded by $(\left\|A+B\widehat{C}_{TR}P_{V}\right\|^{2}+3)t$ provided that the penalty parameter $\lambda$ is large enough. This is related to the intrinsic stability constant associated with the dynamics, underlying manifold. If we assume that two terms $\left\|A+B\widehat{C}_{OLS}P_{V}\right\|,\left\|A+B\widehat{C}_{TR}P_{V}\right\|$ are close to each other, this suggests that our algorithm may have a slower error accumulation rate for large $t$. As the training data lies on a subspace $V$ of the whole state space, we do not have any bound on the behavior of $\widehat{C}_{OLS}$ outside this subspace. For example, the orthogonal complement $V^{\perp}$ may contain an eigenspace of $B\widehat{C}_{OLS}$ with a large eigenvalue. If this is the case, $\left\|B\widehat{C}_{OLS}\right\|$ may be significantly larger than $\left\|A+B\widehat{C}_{TR}P_{V}\right\|$, and the improvement of our algorithm may be more promising. In supplementary material LABEL:fig:linear-cmp, we provide a toy example which illustrates the effectiveness of our method comparing to OLS.

### v2

@@ 3.2 Algorithm in the general form
0:  $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$, resolved model, penalty strength $\lambda$.
1:  Learn a parameterized model which encodes the structure of training data, i.e. $F_{\eta}(\mathbf{u})\geq 0,F_{\eta}(\mathbf{u}_{k})=0.$
@@
6:     Calculate the state variable after one-step iteration, i.e. $\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+\Delta tL(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k},t_{k})$.
7:     Form the loss $l(\theta)=\mathbb{E}\left[\left\|\mathbf{y}_{k}-\widehat{\mathbf{y}}_{k}\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{k})\right)^{T}L(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k},t_{k})\right)^{2}\right]$.
@@
Compared to the literature on distributionally robust optimization (DRO)  [44] and numerical analysis of dynamical systems  [56, 55], our work takes advantages from both. First, the tangent-space regularized estimator belongs to the family of DRO in the sense that it minimizes the data-driven module of the unresolved part over some perturbations of the data distribution. However, unlike the original DRO which considers all possible perturbations of the data distribution in its neighborhood under some metric, e.g. Wasserstein metric [4], our regularization focus on those perturbations caused by the resolved components of the dynamics. These facilitate us to take advantage of the partial knowledge of the dynamics. As a result, the regularizer we proposed is more problem-specific. Moreover, from the perspective of DRO, our method takes into consideration the behavior of estimator under certain perturbation normal to the data manifold, which is much more tractable than guarding against arbitrary perturbations in distribution space. In Section 4, we will implement other benchmarks with general regularizations and compare with our algorithm to illustrate the benefits of combining resolved model information into regularization.
Furthermore, we can understand our method through the lens of the stability analysis of numerical methods for differential equations. The classical result that the least squares estimator converges to the ground truth in the limit of large datasets can be understood as a consistency statement. Here, we show that this is insufficient, and to ensure convergence of the simulated trajectories one needs to also promote stability, and the regularizer proposed by us serves precisely this role. This regularization approach should be contrasted with recent works [62, 6, 29, 16] on learning dynamical models from data which builds stability by specifying the model architecture.
@@ 3.3 Interpretation and analysis of the algorithm
Let us now quantify in the linear case, the gains of our algorithm over the ordinary least squares estimator. We first calculate the exact formula for two estimators. Due to the low-rank structure, the OLS algorithm allows infinitely many solutions. Throughout this section, we make the common choice of the minimum 2-norm solution. Let us first recall two important choices considered in this section will be the OLS estimator and tangent-space regularized estimator, i.e.
(23) $C_{\mathrm{OLS}}$ $:=\arg\min_{C}\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{y}_{i}-C\mathbf{u}_{i}\right\|_{2}^{2},$
$C_{\mathrm{TR}}$ $:=\arg\min_{C}\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\mathbf{y}_{i}-C\mathbf{u}_{i}\right\|_{2}^{2}+\lambda\left\|\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{u}_{i}\right\|_{2}^{2}\right).$
The next proposition provides the explicit formula for them.
@@
Consider the linear dynamics Eq. 12, the training data of the variable $\mathbf{u}$ is arranged into a data matrix $\mathbf{U}\in\mathbb{R}^{m\times N}$. Moreover, we assume the column space of $\mathbf{U}$ is contained in a subspace $V\subset\mathbb{R}^{m}$ with associated projection operator $\operatorname{P}_{V}$. Then, the OLS estimator and tangent-space regularized estimator for the unresolved component $Y=C^{*}\mathbf{U}+\epsilon$ have the following form:
(24) $\widehat{C}_{\mathrm{OLS}}=$ $\ C^{*}\operatorname{P}_{V}+\epsilon\mathbf{U}^{\dagger},$
@@
$\widehat{C}_{\mathrm{TR}}=$ $\ (\mathbf{I}+\lambda B^{T}\operatorname{P}_{V^{\perp}}B)^{-1}(C^{*}\operatorname{P}_{V}+\epsilon\mathbf{U}^{\dagger}-\lambda B^{T}\operatorname{P}_{V^{\perp}}AP_{V}).$
@@
Specifically, in the noiseless scenarios, $\widehat{C}_{\mathrm{OLS}}$ recovers $C^{*}$ on the subspace $V$ but vanishes on its orthogonal complement. To interpret our estimator, let us consider a simpler case with $B=\mathbf{I},A=\mathbf{0}$. This corresponds to the dynamic $\partial_{t}\mathbf{u}=C^{*}\mathbf{u}$, and the regularized estimator is given by
@@
(25) $\widehat{C}_{\mathrm{TR}}=(\mathbf{I}+\lambda\operatorname{P}_{V^{\perp}})^{-1}(C^{*}\operatorname{P}_{V}+\epsilon\mathbf{U}^{\dagger})=(\mathbf{I}+\lambda\operatorname{P}_{V^{\perp}})^{-1}\widehat{C}_{\mathrm{OLS}}.$
@@
Therefore, our estimator performs a weighted least squares regression which penalizes the direction perpendicular to the data subspace $V$, controlled by $\lambda$. As the error along this normal direction is the main cause of the distribution shift, it is sensible to reduce it more than the error along the tangent direction, which causes deviations inside the manifold. This shows that the tangent-space regularized estimator does mitigate distribution shift.
Proof 3.2 (Proof of Proposition 3.1).
We solve the linear regression problem of the unresolved component analytically. Recall that the solution for the $\min_{C}\left\|Y-CU\right\|_{\rm F}^{2}$ is given by $\widehat{C}=YU^{\dagger}$ where $U^{\dagger}$ is the pseudo-inverse of the matrix $U$. Consequently, assume $Y$ is noiseless, i.e. $Y=C^{*}U$, we have $\widehat{C}=C^{*}UU^{\dagger}$. Let us assume the number of data samples is greater than the dimension of the state, i.e. $U\in\mathbb{R}^{n\times N},n<N$. Then, if $U$ is full rank, i.e. $\operatorname{rank}U=n$, the product $UU^{\dagger}$ is exactly the identity matrix and $\widehat{C}$ thereby recovers the $C^{*}$ exactly. However, in our setting, the data is supported on a low dimensional subspace and the data matrix is rank-deficient. Therefore, $UU^{\dagger}=\operatorname{P}_{V}$, where $V$ is the column space of the data matrix $U$ and also the support set of it. We have $\widehat{C}=C^{*}\operatorname{P}_{V}$ in this case. Similarly, in the noisy case $Y=C^{*}U+\epsilon$, we have
(26) $\widehat{C}_{\mathrm{OLS}}=C^{*}\operatorname{P}_{V}+\epsilon U^{\dagger}.$
This is the OLS estimator of the unresolved component.
Next, we calculate the formula for tangent-space regularized estimator. The loss function can be reduced to
(27) $\ \left\|Y-C\mathbf{U}\right\|_{\rm F}^{2}+\lambda\left\|\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{U}\right\|_{\rm F}^{2}$
$=$ $\ \operatorname{Tr}[(Y-C\mathbf{U})^{T}(Y-C\mathbf{U})]+\lambda\operatorname{Tr}[(\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{U})^{T}(\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{U})]$
$=$ $\ \operatorname{Tr}[(Y-C\mathbf{U})^{T}(Y-C\mathbf{U})]+\lambda\operatorname{Tr}[\mathbf{U}^{T}(A+BC)^{T}\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{U}],$
where we use the fact $\operatorname{P}_{V^{\perp}}^{T}\operatorname{P}_{V^{\perp}}=\operatorname{P}_{V^{\perp}}$ since $\operatorname{P}_{V^{\perp}}$ is a projection. Consequently, the first-order condition is given by
(28) $C\mathbf{U}\mathbf{U}^{T}-Y\mathbf{U}^{T}+\lambda B^{T}\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{U}\mathbf{U}^{T})=0.$
We finally derive our estimator as
(29) $\widehat{C}_{\mathrm{TR}}=$ $\ (\mathbf{I}+\lambda B^{T}\operatorname{P}_{V^{\perp}}B)(Y\mathbf{U}^{T}-\lambda B^{T}\operatorname{P}_{V^{\perp}}A\mathbf{U}\mathbf{U}^{T})(\mathbf{U}\mathbf{U}^{T})^{\dagger}$
$=$ $\ (\mathbf{I}+\lambda B^{T}\operatorname{P}_{V^{\perp}}B)^{-1}(C^{*}\operatorname{P}_{V}+\epsilon U^{\dagger}-\lambda B^{T}\operatorname{P}_{V^{\perp}}AP_{V}).$
Remark 3.3.
Unlike classical $L^{1}$ or $L^{2}$-regularization, the tangent-space regularization does not have the issue of bias-variance trade-off, as both the least square and the regularization terms admit a common minimizer $C^{*}$. This is easily observed by plugging in the ground truth $C^{*}$ to the formula of the regularizer, giving
(30) $\left\|\operatorname{P}_{V^{\perp}}(A+BC^{*})\mathbf{u}_{i}\right\|_{2}^{2}=0,$
since $\mathbf{u}_{i}$ lies in the invariant subspace $V$ of the generator of the dynamics $A+BC^{*}$. Therefore, increasing $\lambda$ will mitigate the effect of distribution shift, while at the same time it does not deteriorate the performance of the least squares problem.
@@
(31) $l_{TR}(\widehat{C})=\mathbb{E}\left(\left\|(C^{*}-\widehat{C})\mathbf{u}\right\|_{2}^{2}+\lambda\left\|\operatorname{P}_{V^{\perp}}(A+B\widehat{C})\mathbf{u}\right\|_{2}^{2}\right).$
@@
[…]
where $a\vee b:=\max(a,b).$ In all the discussions that follow, we use $\mathrm{eig}_{\max}(A)$ to denote the greatest real part of the eigenvalues of $A$, i.e.
(33) $\mathrm{eig}_{\max}(A)=\max\{\Re(s):\exists v\neq\mathbf{0},Av=sv\}.$
As there are several assumptions for the theorem, we state it here for clarity. {assumption} We assume both the OLS estimator and $\mathrm{TR}$ estimator $\widehat{C}_{\mathrm{OLS}}$, $\widehat{C}_{TR}$ are optimized to have error bounded by $\delta$ so that $l_{\mathrm{OLS}}(\widehat{C}_{\mathrm{OLS}}),l_{TR}(\widehat{C}_{TR})<\delta$ in Eq. 14 and Eq. 31. We assume also that $\left\|\operatorname{P}_{V^{\perp}}(A+B\widehat{C}_{\mathrm{TR}})\operatorname{P}_{V}\right\|_{2}$ $<\sqrt{\delta/\lambda}$. Without the second assumption, the error bounds in Theorem 3.4 hold with high probability via the observation
(34) $\operatorname{P}_{V^{\perp}}(A+B\widehat{C}_{\mathrm{TR}})\operatorname{P}_{V}\mathbf{u}=\operatorname{P}_{V^{\perp}}B(\mathbf{I}+\lambda B^{T}\operatorname{P}_{V^{\perp}}B)^{-1}\epsilon\mathbf{U}^{\dagger},$
and classical theory of linear regression [19] if the training data has variance $\Sigma\geq c\mathbf{I},c>0$. In Section 3.3, we include this assumption to simplify the proof.
Moreover, we define several notations to simplify the derivation. Let $P_{1},P_{2},P_{3}$ be matrices which transform $A+B\widehat{C}$, $(A+B\widehat{C})\operatorname{P}_{V}$, and $\operatorname{P}_{V^{\perp}}(A+B\widehat{C})$ into their Jordan normal forms respectively and
(35) $c_{1}$ $=\operatorname{cond}(P_{1}),$ $c_{2}=\operatorname{cond}(P_{2}),$ $c_{3}=\operatorname{cond}(P_{3}),$
$e_{1}$ $=\mathrm{eig}_{\max}(A+B\widehat{C}),$ $e_{2}=\mathrm{eig}_{\max}((A+B\widehat{C})\operatorname{P}_{V}),$ $e_{3}=\mathrm{eig}_{\max}(\operatorname{P}_{V^{\perp}}(A+B\widehat{C})),$
where $\operatorname{cond}(P)$ denotes the condition number of matrix $P$.
@@
Under the same setting as Proposition 3.1 and Section 3.3, suppose the true trajectory is simulated from an initial condition $\mathbf{u}(0)\in V$ which follows the same distribution of the training data. Then, the errors of OLS and our algorithm is bounded respectively by
@@
(36) $\ \mathbb{E}\left\|\widehat{\mathbf{u}}_{\mathrm{OLS}}(T)-\mathbf{u}(T)\right\|_{2}\leq c_{1}\sqrt{\delta}\left\|B\right\|_{2}Q_{m}(e_{1},T),$
@@
$\ \mathbb{E}\left\|\widehat{\mathbf{u}}_{\mathrm{TR}}(T)-\mathbf{u}(T)\right\|_{2}\leq c_{2}\sqrt{\delta}\Big{(}\left\|B\right\|_{2}Q_{m}(e_{2},T)$
$\qquad+\frac{9m^{4}c_{3}}{\sqrt{\lambda}}\left(1+3m^{2}c_{1}\sqrt{\delta}\left\|B\right\|_{2}\right)\left\|A+B\widehat{C}\right\|_{2}(1\vee T^{3m})(1\vee e^{e_{1}T})\Big{)}.$
Generally speaking, the exponential factor $e_{1},e_{2}$ in $Q_{m}$ may make both bounds loose, especially for large $T$. This is inevitable as our bounds deal with worst-case scenarios and exponential growth is common in unstable dynamics. If otherwise the dynamic is stable, $Q_{m}$ will decay exponentially and the bound becomes more effective.
Comparing the two bounds, the first difference is the exponents appearing in the $Q$-function, i.e. in OLS it is $e_{1}=\mathrm{eig}_{\max}(A+B\widehat{C})$ while in our algorithm it is $e_{2}=\mathrm{eig}_{\max}((A+B\widehat{C})\operatorname{P}_{V})$. This is consistent with our expectation as the data-driven surrogate can only be trusted in the subspace $V$. Therefore, if the ground truth dynamics is unstable in the orthogonal complement $V^{\perp}$ while stable in $V$, meaning that $e_{1}>0>e_{2}$ the error of OLS will be dominated by the error made in $V^{\perp}$. For tangent-space regularization, the better exponent $e_{2}$ clearly mitigates this. This is also observed quantitatively in Fig. 3: comparing the error propagations of the $\mathrm{TR}$ with $\lambda=10^{7}$ with OLS, the rates are very different at the beginning with $\mathrm{TR}$ close to $e_{2}$ while OLS close to $e_{1}$. However, we need to be careful with the additional terms appearing in the second bound that may have a big $e_{1}$ exponent which may lead to the blow-up of the error. However, it also contains an extra factor $\sqrt{\frac{1}{\lambda}}$ and will be much smaller than the counterpart appearing in OLS bound as long as the penalty $\lambda$ is large enough. This is justified from Fig. 3, where we observe $\mathrm{TR}$ trajectories with increasing $\lambda$s spend more time with a slower rate $e_{2}$, and then saturate to the $e_{1}$-rate.
Figure 3: We compare the simulation error of various estimators in a synthetic linear dynamic where the data subspace $V(V^{\perp})$ is the stable (unstable) manifold. This quantitatively supports our theory from three perspectives: 1). The trajectory error of the OLS simulator coincides with the rate $e_{1}$ over the unstable manifold; 2). The trajectory error of our algorithm has two stages, i.e. $e_{2}$-rate at the beginning and $e_{1}$ later, which correspond exactly to two parts in Eq. 36. 3). As $\lambda\rightarrow\infty$, the length of the first stage increases, meaning a smaller simulation error.
Proof 3.5 (Proof of Theorem 3.4).
[…]
(37) $\frac{d}{dt}\left(\widehat{\mathbf{u}}-\mathbf{u}\right)=$ $\ \left(A+B\widehat{C}\right)\left(\widehat{\mathbf{u}}-\mathbf{u}\right)+B\left(\widehat{C}-C^{*}\right)\mathbf{u},$
which can be solved exactly to obtain
[…]
Therefore, using the bound (proof in supplementary materials) that for $M\in\mathbb{R}^{m\times m}$, $\left\|e^{Mt}\right\|\leq\operatorname{cond}(P)m^{2}(2+t^{m-1})e^{\mathrm{eig}_{\max}(M)t}$, we obtain
[…]
$\leq$ $\ c_{1}\sqrt{\delta}\left\|B\right\|Q_{m}(e_{1},T)\lesssim\sqrt{\delta}Q_{m}(e_{1},T).$
Notice that this bound also holds for $\mathbf{u}_{\mathrm{TR}}$ as it only requires $\mathbb{E}\left\|(C_{\mathrm{TR}}-C^{*})\mathbf{u}(t)\right\|<\delta$ which is guaranteed by our assumption for $l_{\mathrm{TR}}$. Next, for tangent-space regularization, we need to take into account the existence of the low-dimensional subspace $V$ and the projection operator $\operatorname{P}_{V}$. Recall the decomposition of the error propagated dynamics
(40) $\frac{d}{dt}(\widehat{\mathbf{u}}-\mathbf{u})=(A+B\widehat{C})\operatorname{P}_{V}(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}-C^{*})\mathbf{u}+(A+B\widehat{C})(\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}}),$
and the integration formula for the propagation error Eq. 17. We need to control the term $\operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}(t)=\widehat{\mathbf{u}}(t)-\operatorname{P}_{V}\widehat{\mathbf{u}}(t)$ originating from the subspace decomposition.
(41) $\frac{d}{dt}\operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}=$ $\ \operatorname{P}_{V^{\perp}}(A+B\widehat{C})\operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}+\operatorname{P}_{V^{\perp}}(A+B\widehat{C})\operatorname{P}_{V}\left(\widehat{\mathbf{u}}-\mathbf{u}\right)+\operatorname{P}_{V^{\perp}}(A+B\widehat{C})\mathbf{u},$
whose solution can be organized as
(42) $\ \operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}(T)$
[…]
Now, using the Section 3.3, one has
(43) $\ \mathbb{E}\left\|\operatorname{P}_{V^{\perp}}\left(\widehat{\mathbf{u}}_{\mathrm{TR}}(T)\right)\right\|$
$\leq$ $\ c_{3}\sqrt{\frac{\delta}{\lambda}}\left(Q_{m}(e_{3},T)+m^{2}\int_{0}^{T}e^{(T-t)e_{3}}(2+(T-t)^{m-1})\mathbb{E}\left\|\widehat{\mathbf{u}}_{\mathrm{TR}}(t)-\mathbf{u}(t)\right\|\right)$
$\leq$ $\ c_{3}\sqrt{\frac{\delta}{\lambda}}\left(3m^{2}(1\vee T^{m})(1\vee e^{e_{3}T})+9m^{4}c_{1}\sqrt{\delta}\left\|B\right\|(1\vee T^{2m})(1\vee e^{e_{1}T})\right)$
$\leq$ $\ 3m^{2}c_{3}\sqrt{\frac{\delta}{\lambda}}\left(1+3m^{2}c_{1}\sqrt{\delta}\left\|B\right\|\right)(1\vee T^{2m})(1\vee e^{e_{1}T}).$
In the third line, we use the bound Eq. 39. Now we plug in the bounds of $\operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}_{\mathrm{TR}}(T)$ to the error propagation equation and obtain the following bound
(44) $\ \mathbb{E}\left\|\widehat{\mathbf{u}}_{\mathrm{TR}}(T)-\mathbf{u}(T)\right\|\leq c_{2}\sqrt{\delta}\Big{(}\left\|B\right\|Q_{m}(e_{2},T)$
$\qquad+\frac{9m^{4}c_{3}}{\sqrt{\lambda}}\left(1+3m^{2}c_{1}\sqrt{\delta}\left\|B\right\|\right)\left\|A+B\widehat{C}\right\|(1\vee T^{3m})(1\vee e^{e_{1}T})\Big{)}.$
Remark 3.6.
Recall our definition of the distribution shift in Eq. 10. In the proof, assuming $e_{1}>0>e_{2}$, we obtain the following estimates of the simulation error of OLS and our algorithm:
(45) $\mathbb{E}\left\|\operatorname{P}_{V^{\perp}}\left(\widehat{\mathbf{u}}_{\mathrm{OLS}}(T)\right)\right\|$ $=O((1\vee T^{m})(1\vee e^{e_{1}T})),$
$\mathbb{E}\left\|\operatorname{P}_{V^{\perp}}\left(\widehat{\mathbf{u}}_{\mathrm{TR}}(T)\right)\right\|$ $=O(\frac{(1\vee T^{3m})(1\vee e^{e_{1}T})}{\sqrt{\lambda}}).$
Therefore, if we define
(46) $d(\rho,\widehat{\rho})=\Big{|}\mathbb{E}\left\|\operatorname{P}_{V^{\perp}}\mathbf{u}\right\|-\mathbb{E}\left\|\operatorname{P}_{V^{\perp}}\widehat{\mathbf{u}}\right\|\Big{|},\quad\mathbf{u}\sim\rho,\widehat{\mathbf{u}}\sim\widehat{\rho},$
we conclude that the OLS algorithm suffers from distribution shift as $\operatorname{P}_{V^{\perp}}\mathbf{u}=\mathbf{0}$ while our algorithm can mitigate this issue by choosing a large penalty $\lambda$, which is also observed quantitatively in Fig. 3. Although this constant may be exponentially large over $T$ in unstable linear dynamics, our numerical experiments in the next section imply that $\lambda$ needs not to be very large practically.